### v4

@@ 2.1 Optimality conditions
In this subsection, we introduce the first and the second-order optimality condition of problem (P). First, let us recall the definition of subgradient and the first-order optimality condition.
Definition 2.
@@
    $v$ is a regular subgradient of $f$ at $\bar{x}$, written $v\in\hat{\partial}f(\bar{x})$ if
    ^$\hat{\partial}f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$ (2)
@@
    $v$ is a (general) subgradient of $f$ at $\bar{x}$, written $v\in\partial f(\bar{x})$, if there are sequences $x^{v}\rightarrow\bar{x}$ and $v^{v}\in\hat{\partial}f(x^{v})$ with $x^{v}\rightarrow v$.
@@
$\partial\psi(x)=\hat{\partial}\psi(x)=\partial\sum\limits_{i=1}^{n}s(x_{i})=\partial s(x_{1})\times...\times\partial s(x_{n})$ (3)
@@
$\partial s(x_{i})=\hat{\partial}s(x_{i})=\begin{cases}\{\text{sign}(x_{i})r^{\prime}(|x_{i}|)\},&i\in\mathcal{I}(x),\\ [-r^{\prime}(|0^{+}|),+r^{\prime}(|0^{+}|)],&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)<\infty,\\ \mathbb{R},&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)=\infty.\end{cases}$ (4)
Proof.
@@
Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$. By [42, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$. On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, which means $r^{\prime}(|0^{+}|)\in\hat{\partial}s(0),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$. It follows from [42, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$. Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
@@
In general, we have shown that (4) is true. It then follows from [42, Proposition 10.5] that (3) is true. ∎
We can appeal to the classical rule of Fermat [42, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a stationary point of (P), then
$0\in\partial F(x^{*})=\nabla f(x^{*})+\lambda\partial\psi(x^{*})$ (5)
@@
Proposition 3.
If $x^{*}$ is a local minimum for (P), then the following holds
$\nabla_{i}f(x^{*})=-\text{sign}(x_{i}^{*})\lambda r^{\prime}(|x^{*}_{i}|),$ $i\in\mathcal{I}^{*},$ (6)
$\nabla_{i}f(x^{*})\in(-\lambda r^{\prime}(|0^{+}|),+\lambda r^{\prime}(|0^{+}|)),$ $i\in\mathcal{J}^{*}.$ (7)
Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point that satisfies the necessary optimality conditions (6) and (7) is referred to as the stationary point.
In the following, we demonstrate that the second-order optimality condition can be characterized by the nonzero components.
Theorem 4.
@@
    (necessary) If $x^{*}$ is a local minimum of $F$, then
    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (8)
@@
    (sufficient) A stationary point $x^{*}$ of $F$ is a strict local minimum if
    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (9)
@@
Proof.
(i) is proved in [43, Theorem 3.3]. We only prove (ii).
Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of nonzero $\{u^{k}\}$ satisfying
$F(x^{*}+u^{k})\leq F(x^{*})\quad{\rm and}\quad\lim\limits_{k\rightarrow\infty}u^{k}=0.$ (10)
Consider $F_{|\mathcal{I}^{*}}$, which is locally smooth around $x^{*}_{\mathcal{I}^{*}}$ and $x^{*}_{\mathcal{I}^{*}}$ is stationary for $F_{|\mathcal{I}^{*}}$. It follows from $u^{k}=u_{\mathcal{I}^{*}}^{k}+u_{\mathcal{J}^{*}}^{k}$ that
$F(x^{*}+u^{k})-F(x^{*})$ (11)
@@
$-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
$=$ $\ f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}^{k}r(|u_{i}^{k}|).$
From the classic optimality condition for smooth function and (9), it is easy to see that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$. Therefore for sufficiently large $k$,
$F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})>0.$ (12)
Moreover, by the Lagrangian mean value theorem, there exists sufficiently small $\epsilon^{k}\in(0,\min_{i}\{|u_{i}^{k}|\})$ such that
$r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$ (13)
The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}\|_{\infty})$, which, together with (13), yields that
@@
$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|,$ (14)
where $\eta^{k}:=\frac{f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})}{\|u_{\mathcal{J}^{*}}^{k}\|}$. By the definition of gradient,
$\lim_{u^{k}\to 0}\frac{f(x^{*}+u_{\mathcal{J}^{*}}^{k})-f(x^{*})-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\lim_{u^{k}\to 0}\eta^{k}-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=0,$
implying $\limsup\limits_{k\to\infty}\eta^{k}=\limsup\limits_{k\to\infty}\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}\leq\|\nabla f(x^{*})\|_{\infty}$. By (7), $\limsup\limits_{k\to\infty}\eta^{k}\leq\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$. Therefore, $\eta^{k}<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$. It follows from (14) that for sufficiently large $k$,
$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|>0,$
contradicting (10). ∎
@@ 2.2 Strict Saddle Points
In this section, we define the strict saddle point of the problem (P). In smooth settings, the strict saddle can be defined as the points with negative minimum eigenvalue of the Hessian matrix, which indicates a descent direction. For our case, Theorem 4 implies that a saddle point can be defined using nonzeros.
Definition 3.
A stationary point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$. If all stationary points of $F(x)$ are strict saddle points or local minimums, $F$ satisfies the strict saddle property.
@@
To show that the strict saddle property is mathematically generic, many works show that this property holds for a full measure of linear perturbation [23, Theorem 2.9], [44, Theorem 2]. Namely, the perturbed function has the strict saddle property for almost all $v\in\mathbb{R}^{n}$. In our case, we also have the same results by considering the following perturbed nonsmooth regularization problem for $v\in\mathbb{R}^{n}$,
$\min\limits_{x\in\mathbb{R}^{n}}F_{v}(x):=F(x)-\langle v,x\rangle=f(x)+\lambda\psi(x)-\langle v,x\rangle.$ (15)
To show the desired result, we need the assistance of the famous Sard’s theorem. Our proof of the strict saddle property mainly follows the proof of [45, Lemma 2.7], which is first proved in [46].
Definition 4.
Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, if the derivative at $x$ is singular (Jacobian is not full rank), then $x$ is called a critical point of $f$.
Theorem 5.
(Sard’s Theorem) Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, and let $C$ be the set of stationary points of $f$ in $\mathcal{M}$. Then $f(C)$ has measure zero in $\mathcal{N}$.
The genericity of the strict saddle property is stated below.
Theorem 6.
For a full Lebesgue measure set of perturbations $v\in\mathbb{R}^{n}$, the perturbation function $F_{v}(x):=F(x)-\langle v,x\rangle$ satisfies
$\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d\neq 0,$ (16)
for all stationary points $x$ of $F_{v}$ with $\mathcal{I}=\mathcal{I}(x)$.
Proof.
@@
be a manifold. Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$ with $\mathcal{I}=\mathcal{I}(s)$. The Jacobian of $v$ at $x$ is then given by $Dv(x)=\begin{bmatrix}\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)&0\\ 0&I\end{bmatrix}$, and $Dv(x)$ has full rank if and only if $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$ has full rank. By Sard’s Theorem,
$\{v(x)\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n},$
which is equivalent to
$\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}(s)|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}.$
@@
has measure 0 in $\mathbb{R}^{n}$. ∎
Center Stable Manifold Theorem
As many other works in smooth setting like [47], we interpret our algorithm as the fixed point iteration of a nonlinear map $T$. Likewise, we define the fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as the unstable fixed points, as described below.
Definition 5.

### v5

@@ 2.1 Optimality conditions
We discuss the first- and second-order optimality conditions of problem (P) in this subsection. First, let us recall the definition of subgradients and the first-order optimality condition.
Definition 2.1.
@@
    $v$ is a regular subgradient of $f$ at $\bar{x}$, if $v\in\hat{\partial}f(\bar{x})$ where
    (1) ^$\hat{\partial}f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$
@@
[…]
@@
(2) $\partial\psi(x)=\hat{\partial}\psi(x)=\partial\sum\limits_{i=1}^{n}s(x_{i})=\partial s(x_{1})\times...\times\partial s(x_{n})$
@@
(3) $\partial s(x_{i})=\hat{\partial}s(x_{i})=\begin{cases}\{\text{sign}(x_{i})r^{\prime}(|x_{i}|)\},&i\in\mathcal{I}(x),\\ [-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)],&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)<\infty,\\ \mathbb{R},&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)=\infty.\end{cases}$
Proof 2.3.
@@
Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$. By [34, Exercise 8.31], we have $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$. On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ holds for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, which means $r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$ and $-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$. It follows from [34, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$. Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
@@
In general, we have shown that (3) is true. It then follows from [34, Proposition 10.5] that (2) is true.
We can appeal to the classical rule of Fermat [34, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a local minimum of (P), then
(4) $0\in\partial F(x^{*})=\nabla f(x^{*})+\lambda\partial\psi(x^{*})$
@@
Proposition 2.4.
If $x^{*}$ is a local minimum for (P), then the following hold
(5) $\nabla_{i}f(x^{*})=-\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x^{*}_{i}|),$ $i\in\mathcal{I}^{*},$
(6) $\nabla_{i}f(x^{*})\in(-\lambda r^{\prime}(|0^{+}|),+\lambda r^{\prime}(|0^{+}|)),$ $i\in\mathcal{J}^{*}.$
Notice that if $r^{\prime}(|0^{+}|)=\infty$, (6) holds naturally true. Any point that satisfies the necessary optimality conditions (5) and (6) is referred to as a stationary point.
In the following, we demonstrate that the second-order optimality condition can be characterized by nonzero components.
Theorem 2.5.
@@
    (necessity) If $x^{*}$ is a local minimum of $F$, then
    (7) $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$
@@
    (sufficiency) A stationary point $x^{*}$ of $F$ is a strict local minimum if
    (8) $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}}\setminus\{0\},$
@@
Proof 2.6.
(i) is proved in [8, Theorem 3.3]. We only prove (ii).
Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of nonzero vectors $\{u^{k}\}$ satisfying
(9) $F(x^{*}+u^{k})\leq F(x^{*})\quad{\rm and}\quad\lim\limits_{k\rightarrow\infty}u^{k}=0.$
Consider $F_{|\mathcal{I}^{*}}$, which is locally smooth around $x^{*}_{\mathcal{I}^{*}}$ and $x^{*}_{\mathcal{I}^{*}}$ is stationary for $F_{|\mathcal{I}^{*}}$. It follows from $u^{k}=\begin{bmatrix}u_{\mathcal{I}^{*}}^{k};u_{\mathcal{J}^{*}}^{k}\end{bmatrix}$ that
(10) $F(x^{*}+u^{k})-F(x^{*})$
@@
$-f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}])+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
$=$ $\ f(x^{*}+u^{k})-f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}])+F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}^{k}r(|u_{i}^{k}|).$
From the classic optimality condition for a smooth function and (8), it is easy to see that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$. Therefore for sufficiently large $k$,
(11) $F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})>0.$
Moreover, by the Lagrangian mean value theorem,
(12) $r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$
The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}_{\mathcal{J}^{*}}\|_{\infty})$, which, together with (12), yields that
@@
(13) $F(x^{*}+u^{k})-F(x^{*})>(\lambda r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})+\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|,$
where
$\eta^{k}:=\frac{f(x^{*}+u^{k})-f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}])}{\|u_{\mathcal{J}^{*}}^{k}\|}=\nabla f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}]+t[\boldsymbol{0};u_{\mathcal{J}^{*}}^{k}])^{T}[\boldsymbol{0};u_{\mathcal{J}^{*}}^{k})]/\|u_{\mathcal{J}^{*}}^{k}\|,$
and $t\in[0,1]$, implying
$\limsup\limits_{k\to\infty}|\eta^{k}|\leq\limsup\limits_{k\to\infty}\|\nabla f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}]+t[\boldsymbol{0};u_{\mathcal{J}^{*}}^{k}])\|_{\infty}\|\tfrac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}\|_{1}=\|\nabla f(x^{*})\|_{\infty}.$
It follows from $\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$ by (6) that $|\eta^{k}|<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$. It follows from (13) that for sufficiently large $k$,
$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})+\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|>0,$
contradicting (9).
@@ 2.2 Strict Saddle Points
In this subsection, we define the strict saddle point of the problem (P). In smooth settings, the strict saddle point can be defined by virtue of a negative minimum eigenvalue of the corresponding Hessian, indicating a descent direction. For our case, Theorem 2.5 implies that a saddle point can be defined using nonzeros.
Definition 2.7.
A stationary point $x^{*}$ of $F(x)$ is a strict saddle point if
$d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$
for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$. If all stationary points of $F(x)$ are strict saddle points or local minima, then we say that $F$ satisfies the strict saddle property.
@@
To show that the strict saddle property is mathematically generic, many works show that this property holds for a full measure of linear perturbation [15, Theorem 2.9], [4, Theorem 2]. Namely, the perturbed function has the strict saddle property for almost all linear perturbations. In our case, we also have the same results by considering the following linearly perturbed nonsmooth regularization problem for $v\in\mathbb{R}^{n}$,
(14) $\min\limits_{x\in\mathbb{R}^{n}}F_{v}(x):=F(x)-\langle v,x\rangle=f(x)+\lambda\psi(x)-\langle v,x\rangle.$
To show the desired result, we need the assistance of the famous Sard’s theorem [matsumoto2002introduction]. Our proof of the strict saddle property mainly follows the proof of [40, Lemma 2.7], which is first proved in [milnor2025lectures].
Definition 2.8.
Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds. If the derivative at $x$ is singular (Jacobian is not of full rank), then $x$ is called a critical point of $f$.
Theorem 2.9.
(Sard’s Theorem) Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, and $C$ be the set of critical points of $f$ in $\mathcal{M}$. Then $f(C)$ has measure zero in $\mathcal{N}$.
The genericity of the strict saddle property is stated in the following.
Theorem 2.10.
Let $\mathcal{I}$ stand for $\mathcal{I}(x)$. For a full Lebesgue measure set of perturbations $v\in\mathbb{R}^{n}$, the perturbation function $F_{v}(x):=F(x)-\langle v,x\rangle$ satisfies
(15) $\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d\neq 0,$
for all stationary points $x$ of $F_{v}$.
Proof 2.11.
@@
be a manifold. Consider the behavior of $F_{v}$ in $\mathcal{M}(s)$, where $\mathcal{I}=\mathcal{I}(x)=\mathcal{I}(s)$ for all $x\in\mathcal{M}(s)$. Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$. The Jacobian of $v$ at $x$ is then given by $Dv(x)=\begin{bmatrix}\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)&0\\ 0&I\end{bmatrix}$, and $Dv(x)$ has full rank if and only if $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$ has full rank. By Sard’s Theorem, $\{v(x)\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$, which implies its subset
$\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}(s)|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}(s)\}$
has measure 0 in $\mathbb{R}^{n}$.
@@
has measure 0 in $\mathbb{R}^{n}$.
2.3 Center Stable Manifold Theorem
As with many other works in smooth settings like [24], we will interpret our algorithm as the fixed point iteration of a nonlinear map $T$. Likewise, we define fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as unstable fixed points, as described below.
Definition 2.12.